Given the following information, construct the rejection region. Show the solution in


a step-by-step procedure.


H 0 : = 45

H a : < 45

= 40, = 12, n = 32, = 0.01


1
Expert's answer
2022-05-19T19:06:10-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=45"

"H_a:\mu\not=45"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=31" degrees of freedom, and the critical value for a two-tailed test is "t_c =2.744042."The rejection region for this two-tailed test is "R = \{t:|t|> 2.744042\}."

The t-statistic is computed as follows:


"t=\dfrac{40-45}{12/\sqrt{32}}=-2.357"

Since it is observed that "|t| = 2.357<2.744042=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed "df=31" degrees of freedom, "t=-2.357" is "p=0.024919," and since "p=0.024919>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 45, at the "\alpha = 0.01" significance level.